### v2

# Inexact Restoration via random models for unconstrained noisy optimization

Abstract.
We study the Inexact Restoration framework with random models for minimizing functions whose evaluation is subject to errors. We propose a constrained formulation that includes well-known stochastic problems and an algorithm applicable when the evaluation of both the function and its gradient is random and a specified accuracy of such evaluations is guaranteed with sufficiently high probability. The proposed algorithm combines the Inexact Restoration framework with a trust-region methodology based on random first-order models. We analyse the properties of the algorithm and provide the expected number of iterations performed to reach an approximate first-order optimality point. Numerical experiments show that the proposed algorithm compares well with a state-of-the-art competitor.
Key words and phrases:
Trust-region algorithms, random models, Inexact Restoration.
2020 Mathematics Subject Classification:
65K05, 90C30, 90C15
The authors are members of the INdAM Research Group GNCS. The research that led to the present paper was partially supported by INDAM-GNCS through Progetti di Ricerca 2024 and by PNRR - Missione 4 Istruzione e Ricerca - Componente C2 Investimento 1.1, Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN) funded by the European Commission under the NextGeneration EU programme, project “Advanced optimization METhods for automated central veIn Sign detection in multiple sclerosis from magneTic resonAnce imaging (AMETISTA)”, code: P2022J9SNP, MUR D.D. financing decree n. 1379 of 1st September 2023 (CUP E53D23017980001), project “Numerical Optimization with Adaptive Accuracy and Applications to Machine Learning”, code: 2022N3ZNAX MUR D.D. financing decree n. 973 of 30th June 2023 (CUP B53D23012670006), and by Partenariato esteso FAIR “Future Artificial Intelligence Research” SPOKE 1 Human-Centered AI. Obiettivo 4, Project “Mathematical and Physical approaches to innovative Machine Learning technologies (MaPLe)”, code: EP_FAIR_002, CUP B93C23001750006.
1. Introduction
We consider unconstrained optimization problems of the form
(1.1) $\min_{x\in\mathbb{R}^{n}}f(x),$
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a smooth function whose evaluation is inexact. In particular, we are interested in problems where the evaluation of both the objective function and its gradient is random and sufficient accuracy in the estimates can be guaranteed with sufficiently high probability [16, 22]. This class of problems includes the minimization of either functions which are intrinsically noisy or expensive functions which are approximately computed, see, e.g., [10, 11, 15].
Our proposal for the numerical solution of this problem is based on the Inexact Restoration framework for constrained optimization problems proposed in [21]. Inexact Restoration is a two-phase method consisting of a restoration phase and an optimization phase; the restoration phase improves feasibility and does not call for function evaluations, whereas the optimization phase improves the objective function value with respect to the point obtained in the restoration phase, see e.g., [1, 12, 14, 17]. Such a framework has been successfully employed in the minimization of functions whose computation is intrinsically inexact; specifically, the exact evaluation of the function is interpreted as feasibility of a constrained problem and the restoration phase increases the accuracy in the function evaluation [18, 7, 8]. Analogously, Inexact Restoration has been fruitfully applied in finite-sum minimization with inexact functions and derivatives [5, 4, 6].
In order to apply the Inexact Restoration approach, we reformulate (1.1) as
(1.2) $\min_{x\in\mathbb{R}^{n}}\bar{f}(x,y,\omega_{y})\quad\mbox{ subject to }h(y)=0,$
where $\bar{f}(x,y,\omega_{y})$ is an estimate of $f(x)$ with noise level $y\in Y$, $\omega_{y}$ denotes a random variable dependent of $y$ and independent of $x$, and $h:Y\rightarrow\mathbb{R}$ is a nonnegative function such that the lower $h(y)$, the higher the accuracy of the function and gradient estimates is in probability. In the ideal case $h(y)=0$, the function estimate $\bar{f}(x,y,\omega_{y})$ coincides with the exact objective value $f(x)$. We show that our problem setting is viable for well-known optimization problems and then design and analyze a new trust-region algorithm that employs random first-order models for optimization. We borrow ideas from the Inexact Restoration approach and adapt such framework, in order to define an algorithm that is well-defined and theoretically well founded even if the random models are inaccurate at some iterations. We denote our algorithm as irerm (Inexact REstoration with Random Models).
Our work is different from existing works on the Inexact Restoration approach for the minimization of noisy functions [13, 18, 7, 8], where either the noise is controllable in a deterministic way or the noise is stochastic but the probability of obtaining sufficiently accurate estimates of functions and gradients is increasing; by contrast, we assume that the noise is stochastic and the probability of obtaining accurate estimates is prefixed. A stochastic first-order trust-region algorithm with Inexact Restoration has been proposed in [4], but only for finite-sum minimization problems; by contrast, irerm can be applied to any unconstrained differentiable problem of the form (1.1). Our algorithm falls in the class of trust-region procedures with random models, see e.g., the contributions [2, 3, 9, 15, 19]. With respect to the existing literature, irerm is inspired by storm, the trust-region algorithm with random models proposed in [15] and further analysed in [9]. Differently from storm, the probabilistic accuracy requirements on function and gradient estimates are explicitly imposed within irerm using problem (1.2) and a suitable function $h$. Furthermore, the acceptance test in irerm is a stochastic counterpart of the rule typically adopted in the Inexact Restoration framework and involves both $h$ and the noisy function $\bar{f}$, whereas the acceptance test in storm is a stochastic counterpart of the classical trust-region acceptance test and involves only the noisy values of $f$.
We design the new algorithm irerm and show that it shares theoretical properties analogous to those of storm. Particularly, we provide a bound on the expected number of iterations needed to achieve a certain accuracy in the norm of the true gradient. Furthermore, we compare numerically irerm with storm on a collection of least-squares problems and observe that irerm achieves comparable or lower noiseless values of $f$ than storm.
The rest of the paper is organized as follows. In Section 2 we discuss our constrained problem formulation (1.2) and present fields of applications for our approach. In Section 3 we introduce our procedure and in Section 4 we study its theoretical properties and provide the expected number of iterations performed to reach an approximate first-order optimality point for (1.1). Finally, in Section 5 we provide the numerical validation of our new algorithm.
2. Problem formulation
In this section, we discuss the assumptions made on problem (1.1) and its constrained formulation (1.2) and show that they are viable.
For any $x\in\mathbb{R}^{n}$ and level of noise $y\in Y$, let $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ be random evaluations of $f(x)$ and $\nabla f(x)$, respectively, where $\omega_{y}$ denotes a random variable from a probability space with probability measure $\mathbb{P}$, dependent of $y$ and independent of $x$, that models the noise on the function and its gradient. Regarding the constraint in (1.2), the function $h:Y\rightarrow\mathbb{R}$ is assumed to be non-negative. If $h(y)=0$, then $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ coincide with $f(x)$ and $\nabla f(x)$, respectively. If $h(y)$ is non-zero, its value is related to the level of accuracy that $\bar{f}(x,y,\omega_{y})$ and $\nabla_{x}\bar{f}(x,y,\omega_{y})$ can achieve in probability with respect to the corresponding exact evaluations. We summarize these assumptions in the following.
Assumption 2.1.
  • (i)
    There exist a function $h:Y\rightarrow\mathbb{R}$ and $h_{up}\geq 0$ such that $0\leq h(y)\leq h_{up}$ for all $y\in Y$. If $h(y)=0$, then $\bar{f}(x,y,\omega_{y})=f(x)$ for all $x\in\mathbb{R}^{n}$.
  • (ii)
    Given $x\in\mathbb{R}^{n}$, there exists a function $\bar{\mu}:(0,1]\rightarrow\mathbb{R}$, $\bar{\mu}(\alpha)\in[0,\bar{\mu}_{up}]$ for all $\alpha\in(0,1]$ and for some $\bar{\mu}_{up}>0$, such that, if $y\in Y$ satisfies $h(y)\leq\bar{\mu}(\alpha)\rho$, with $\alpha\in(0,1]$, $\rho>0$, then it holds
    (2.1) $\mathbb{P}(|f(x)-\bar{f}(x,y,\omega_{y})|\leq\rho)\geq\alpha,\quad\mathbb{P}(\|\nabla f(x)-\nabla_{x}\bar{f}(x,y,\omega_{y})\|\leq\rho)\geq\alpha.$
  • (iii)
    Functions $f(x)$ and $\bar{f}(x,y)$ are bounded from below, i.e., there exists $f_{low}$ such that
    $f(x)$ $\geq f_{low},\quad\forall\ x\in\mathbb{R}^{n}$
    $f(x,y,\omega_{y})$ $\geq f_{low},\quad\forall\ x\in\mathbb{R}^{n},\ \forall\ y\in Y.$
Our problem setting includes, but it is not limited to, the following three settings of stochastic noise in unconstrained optimization.
First, in finite-sum minimization problems, where the objective function in (1.1) has the form
(2.2) $f(x)=\frac{1}{N}\sum_{i=1}^{N}\phi_{i}(x),$
and $N$ is a large positive integer, approximations of $f$ and $\nabla f$ can be computed by sampling. This amounts to fixing the sample size $y$ as an integer in $Y=\{1,\ldots,N\}$, choosing the sample set of indices $I_{y}\subset\{1,\ldots,N\}$ of cardinality $y$ at random, and computing
(2.3) $\bar{f}(x,y,\omega_{y})=\frac{1}{y}\sum_{i\in I_{y}}\phi_{i}(x),\quad\nabla_{x}\bar{f}(x,y,\omega_{y})=\frac{1}{y}\sum_{i\in I_{y}}\nabla\phi_{i}(x).$
By an abuse of notation, here $\omega_{y}$ is a realization of the noise, represented by the random selection of a subset $I_{y}$ of cardinality $y$ containing the sample indices. Such evaluations provide the exact values $f(x)$ and $\nabla f(x)$ if $y=N$, while they are sufficiently accurate in probability if $y\in\{1,\ldots,N-1\}$ is large enough [23, §6]. Thus, problem (1.1) with (2.2) as objective function can be reformulated as (1.2) with
$h(y)=\begin{cases}\sqrt{\frac{N}{y}},\quad&\text{if }1\leq y\leq N-1\\ 0,\quad&\text{if }y=N.\end{cases}$
Regarding the fulfillment of Assumption 2.1(ii), we assume that the noise in function evaluations is unbiased and we let $\operatorname{Var}_{\omega_{y}}[\bar{f}(x,y,\omega_{y})]\leq V<\infty$. Then, given $\rho>0$, we choose $y$ such that
(2.4) $h(y)\leq\sqrt{\frac{N(1-\alpha)}{V}}\rho.$
If $y=N$, such condition is satisfied for any $\rho>0$ being $h(N)=0$, and the probabilistic inequalities in (2.1) trivially hold. If $y<N$, condition (2.4) is equivalent to
(2.5) $y\geq\frac{V}{(1-\alpha)\rho^{2}}.$
and $\bar{f}(x,y,\omega_{y})$ satisfies (2.1), see [23, §6]. The same reasoning applies to the evaluation of $\nabla_{x}\bar{f}(x,y,\omega_{y})$ letting $V$ be the upper bound for the variance $\operatorname{Var}_{\omega_{y}}[\nabla_{x}\bar{f}(x,y,\omega_{y})]$. Thus, Assumption 2.1(ii) holds with $\bar{\mu}(\alpha)=\sqrt{N(1-\alpha)/V}$ and $\bar{\mu}_{up}=\sqrt{N/V}$.
Second, consider the case where the objective function in (1.1) is the expectation of a stochastic function that depends on a random variable $\xi$, i.e.,
(2.6) $f(x)=\mathbb{E}_{\xi}[f(x,\xi)],\,\quad\forall\ x\in\mathbb{R}^{n}.$
Sample averaging approximations for reducing the variance on the evaluations can be used, and we set
(2.7) $\bar{f}(x,y,\omega_{y})$ $=$ $\begin{cases}f(x),\quad&\text{if }y=0\\ \displaystyle\frac{1}{p(y)}\sum_{i=1}^{p(y)}f(x,\xi_{i}),\quad&\text{if }y\in(0,1],\end{cases}$
(2.8) $\nabla_{x}\bar{f}(x,y,\omega_{y})$ $=$ $\begin{cases}\nabla f(x),\quad&\text{if }y=0\\ \displaystyle\frac{1}{p(y)}\sum_{i=1}^{p(y)}\nabla_{x}f(x,\xi_{i}),\quad&\text{if }y\in(0,1],\end{cases}$
where the symbol $p(y)=\lceil 1/y\rceil\in\mathbb{N}$ represents the number of samples for a given $y\in(0,1]$ and $\omega_{y}$ represents the i.i.d. noise realizations of $\xi$, $i=1,\ldots,p(y)$. In this case, sufficient accuracy in probability is attained in the approximations $\bar{f}(x,y),\nabla_{x}\bar{f}(x,y)$ by choosing $y$ small enough [23, §6] and problem (1.1) can be cast into (1.2), for instance by setting $Y=[0,1]$ and $h(y)=\sqrt{y}$, $y\in Y$.
As for Assumption 2.1(ii), we let $\operatorname{Var}_{\omega_{y}}[\bar{f}(x,y,\omega_{y})]\leq V<\infty$ for all $x\in\mathbb{R}^{n}$ and suppose that the noise is unbiased. Given $\rho>0$, choose $y\in Y$ such that
(2.9) $h(y)\leq\sqrt{\frac{1-\alpha}{V}}{\rho}.$
If $y=0$, then the above condition is a satisfied, and the first condition in (2.1) trivially holds, as $\bar{f}(x,y,\omega_{y})=f(x)$ according to (2.7). Otherwise, condition (2.9) can be rewritten as $\frac{1}{y}\geq\frac{V}{(1-\alpha)\rho^{2}}$, which implies
(2.10) $p(y)\geq\frac{V}{(1-\alpha)\rho^{2}}.$
The estimator $\bar{f}(x,y)$ in (2.8) satisfies the first probabilistic condition in (2.1) if the sample size $p(y)$ satisfies (2.10), see [23, §6]. An analogous result holds for the second probabilistic condition in (2.1), under the assumption that the noise in the gradient computation is unbiased, and the variance of the gradient estimate is upper bounded by $V$. Thus, Assumption 2.1(ii) holds with $\bar{\mu}(\alpha)=\sqrt{(1-\alpha)/V}$, $\bar{\mu}_{up}=\sqrt{1/V}$.
Finally, problem (1.2) includes the least-squares problem
$f(x)=\frac{1}{2}\|F(x)\|_{2}^{2},$
associated to a nonlinear system $F(x)=0$ where both $F$ and its Jacobian are computed via Monte Carlo simulations, see e.g., [24]. Evaluations are made by using a number of trials as in (2.7) and (2.8) and enlarging the trial set increases the accuracy as in the case discussed above.
3. An inexact restoration trust region method with random models
In the following, we introduce our algorithm named irerm (Inexact REstoration with Random Models). irerm is a stochastic variant of the classical trust-region algorithm, where suitable approximations of the objective function and its gradient are employed, and whose acceptance criterion relies on the decrease of a convex combination of the function estimate with the infeasibility measure $h$. Now we refer to the algorithm irerm stated below and provide a detailed description of it.
Algorithm irerm : Inexact REstoration with Random Models
Input: $x_{0}\in\mathbb{R}^{n}$, $y_{0}\in Y$, $\eta_{1}\in(0,1)$, $\eta_{2}>0$, $\theta_{0}\in(0,1)$, $\underline{\theta}\in(0,\theta_{0})$, $r\in(0,1)$, $\mu>0$, $\gamma>1$, $\delta_{\max}>0$, $\delta_{0}\in(0,\delta_{\max}]$.
  • 0.
    Set $k=0$.
  • 1.
    Accuracy variables
  • Choose $\widetilde{y}_{k+1}$, $y_{k+1}^{t}$ and $y_{k+1}^{g}$ in $Y$ satisfying
    (3.1) $h(\widetilde{y}_{k+1})\leq rh(y_{k}),\,\,h({y}^{t}_{k+1})\leq\mu\min\{\delta_{k}^{2},h(y_{k})\},\,\,h(y_{k+1}^{g})\leq\mu\delta_{k}.$
  • 2.
    Trust-region model
  • Compute $\bar{f}_{k}^{\dagger}=\bar{f}(x_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{\dagger})$ , $g_{k}=\nabla_{x}\bar{f}(x_{k},y_{k+1}^{g},\omega_{y_{k+1}^{g}})$, set
    $p_{k}=-\delta_{k}\frac{g_{k}}{\|g_{k}\|},$
    and build
    $m_{k}(p)=\bar{f}_{k}^{\dagger}+g_{k}^{T}p.$
  • 3.
    Penalty parameter
  • Compute $\bar{f}_{k}^{*}=\bar{f}(x_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{*})$ and evaluate $\operatorname{Pred}_{k}(\theta_{k})$ as in (3.5).
  • Compute the penalty parameter $\theta_{k+1}^{t}$ as
    (3.2) $\theta_{k+1}^{t}=\left\{\begin{array}[]{ll}\theta_{k},&\mbox{if}\ \operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|\\ \displaystyle\frac{h(y_{k})-h(\widetilde{y}_{k+1})}{\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}+h(y_{k})-h(\widetilde{y}_{k+1})},&\mbox{otherwise}.\end{array}\right.$
  • 4.
    Acceptance test
  • Compute $\bar{f}_{k}^{p}=\bar{f}(x_{k}+p_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{p})$, evaluate $\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1}^{t})$ in (3.7) and $\operatorname{Pred}_{k}(\theta_{k+1}^{t})$ as in (3.5).
  • If $\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1}^{t})\geq\eta_{1}\operatorname{Pred}_{k}(\theta_{k+1}^{t})$, $\|g_{k}\|\geq\eta_{2}\delta_{k}$ , $\theta_{k+1}^{t}\geq\underline{\theta}$ (success), set
    (3.3) $x_{k+1}=x_{k}+p_{k},\quad y_{k+1}=y_{k+1}^{t},\quad\delta_{k+1}=\min\{\gamma\delta_{k},\delta_{\max}\},\quad\theta_{k+1}=\theta_{k+1}^{t}.$
    Else (unsuccess) set
    (3.4) $x_{k+1}=x_{k},\quad y_{k+1}=y_{k},\quad\delta_{k+1}=\delta_{k}/\gamma,\quad\theta_{k+1}=\theta_{k}.$
  • 5.
    Set $k=k+1$ and go to Step 1.
At each iteration $k\geq 0$, we have at our disposal the iterate $x_{k}\in\mathbb{R}^{n}$, the accuracy variable $y_{k}\in Y$, the trust-region radius $\delta_{k}>0$, and the penalty parameter $\theta_{k}\in(0,1)$.
At Step 1, the accuracy variables $\widetilde{y}_{k+1}\in Y$, $y_{k+1}^{t}\in Y$, and $y_{k+1}^{g}\in Y$ are selected according to condition (3.1). The variable $\widetilde{y}_{k+1}$ is chosen so that the value $h(\widetilde{y}_{k+1})$ is smaller than a fraction of the current value $h(y_{k})$; the variable $y_{k+1}^{t}$ is selected so that $h(y_{k+1}^{t})$ is upper bounded by a quantity proportional to the minimum value between $h(y_{k})$ and the squared trust-region radius $\delta_{k}^{2}$; the variable $y_{k+1}^{g}$ is such that $h(y_{k+1}^{g})$ is upper bounded by a quantity proportional to $\delta_{k}$. As shown in the next section, condition (3.1) is crucial to provide sufficiently accurate function and derivative approximations in probability. We observe that the variable $\widetilde{y}_{k+1}$ improves feasibility and appears both in the computation of the parameter $\theta_{k+1}^{t}$ (Step 3) and in the evaluation of $\operatorname{Pred}_{k}(\theta_{k})$ and $\operatorname{Pred}_{k}(\theta_{k+1}^{t})$ (Steps 3-4); however, it is not actually used to compute function or gradient estimates. On the other hand, the variables $y_{k+1}^{t}$ and $y_{k+1}^{g}$ are employed for computing three function estimates and one gradient estimate, respectively, but do not necessarily enforce a decrease of the infeasibility measure.
At Step 2 the function estimate $\bar{f}_{k}^{\dagger}=\bar{f}(x_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{\dagger})$ and the gradient estimate $g_{k}=\nabla_{x}\bar{f}(x_{k},y_{k},\omega_{y_{k+1}}^{g})$ are computed for noise realizations $\omega_{y_{k+1}^{t}}^{\dagger}$ and $\omega_{y_{k+1}}^{g}$, respectively. Then the linear model $m_{k}(p)=\bar{f}_{k}^{\dagger}+g_{k}^{T}p$ is formed, and finally the search direction $p_{k}$ is computed by minimizing the model over the ball of center zero and radius $\delta_{k}$, equivalently
$p_{k}=\underset{\|p\|\leq\delta_{k}}{\operatorname{argmin}}\ m_{k}(p)=-\delta_{k}\frac{g_{k}}{\|g_{k}\|}.$
At Step 3, we compute the trial penalty parameter $\theta_{k+1}^{t}$ by using the predicted reduction $\operatorname{Pred}_{k}(\theta)$ defined as
(3.5) $\operatorname{Pred}_{k}{}(\theta)=\theta(\bar{f}_{k}^{*}-m_{k}(p_{k}))+(1-\theta)(h(y_{k})-h(\widetilde{y}_{k+1})),$
where $\theta\in(0,1)$ and $\bar{f}^{*}_{k}=\bar{f}(x_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{*})$ is a further function estimate. More precisely, the parameter $\theta_{k+1}^{t}$ is computed so that the following condition holds
(3.6) $\operatorname{Pred}_{k}(\theta)\geq{\theta}\delta_{k}\|g_{k}\|.$
If condition (3.6) is satisfied with $\theta=\theta_{k}$, then we set ${\theta_{k+1}^{t}}=\theta_{k}$. Otherwise $\theta_{k+1}^{t}$ is computed as the largest value for which inequality (3.6) holds and its value given in (3.2) is derived in Lemma 4.2.
At Step 4, we accept or reject the trial point $x_{k}+p_{k}$. Given a function estimate $\bar{f}_{k}^{p}=\bar{f}(x_{k}+p_{k},y_{k+1}^{t},\omega_{y_{k+1}^{t}}^{p})$, we define the actual reduction at the trial point $x_{k}+p_{k}$ as
(3.7) $\operatorname{Ared}_{k}(x_{k}+p_{k},\theta)=\theta(\bar{f}_{k}^{*}-\bar{f}_{k}^{p})+(1-\theta)(h(y_{k})-h(y^{t}_{k+1})).$
Then, the iteration is declared successful, i.e., the point $x_{k}+p_{k}$ is accepted, whenever the following three conditions hold
(3.8) $\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1}^{t})$ $\geq\eta_{1}\operatorname{Pred}_{k}(\theta_{k+1}^{t}),$
(3.9) $\|g_{k}\|$ $\geq\eta_{2}\delta_{k},$
(3.10) $\theta_{k+1}^{t}$ $\geq\underline{\theta}.$
Condition (3.8) mimics the classical acceptance criterion of standard trust-region methods, whereas (3.9)-(3.10) are technical conditions on the norm of the gradient estimator and the penalty parameter needed for the theoretical analysis of Algorithm irerm. If conditions (3.8)-(3.10) are all satisfied, we update the iterates $x_{k+1},y_{k+1},\delta_{k+1},\theta_{k+1}$ as in (3.3) and proceed to the next iteration, otherwise we retain the iterate $x_{k}$, update the parameters as in (3.4) and proceed to the next iteration.
Condition (3.10) is not standard for deterministic Inexact Restoration approaches and deserves some comments. Such a condition has been introduced in order to ensure that the sequence $\{\theta_{k}\}$ is bounded away from zero, allowing us to avoid certain technical assumptions, employed in the previous literature, that may not be satisfied in our settings. More precisely, in seminal deterministic algorithms such as [7, 8, 13], it is proved that the sequence $\{\theta_{k}\}$ is bounded away from zero by enforcing a condition on the noisy function values along the iterations. Such a condition is problem dependent and hard to impose unless the function $\bar{f}(x,y,\omega)$ is Lipschitz continuous with respect to $y$ and $\omega$, which is hardly verifiable or satisfied if the function evaluations are random, as is in our case. Likewise, in the stochastic Inexact Restoration algorithms [4–6] for finite-sum minimization, we proved that $\{\theta_{k}\}$ is bounded away from zero assuming that every subsampled function evaluation is uniformly bounded from below and above on a set containing the generated sequence $\{x_{k}\}$. For the sake of generality, we preferred to abandon such assumption and enforce condition (3.10) instead.
We conclude this section by comparing irerm with existing trust-region algorithms with random models. First, consider the algorithm storm given in [9, 15]. On the one hand, Step 1 in irerm enforces sufficient accuracy in probability for the function and gradient estimates under a suitable choice of the parameter $r$ (see the upcoming Lemma 4.10). On the other hand, setting the noise level is not explicitly part of storm. Furthermore, the actual reduction (3.7) used in Step 4 is a convex combination of the reduction in the function approximation from $x_{k}$ to $x_{k}+p_{k}$ and the reduction in the function $h$. Thus, irerm may accept iterations where no decrease occurs in function approximations, provided that there has been a sufficient improvement towards the accuracy constraint $h(y)=0$. Such an acceptance rule differs from that in storm, which is solely based on function estimates.
Second, we highlight some important differences between irerm and sirtr, the stochastic trust-region algorithm with Inexact Restoration proposed in [4]. Algorithm sirtr is applicable only to finite-sum minimization problems and its update rule for the accuracy variable is merely based on feasibility restoration, while not ensuring the probabilistic accuracy requirements (2.1) for random functions and gradients. Consequently, it may be necessary to further enlarge the sample size originated by the update rule of sirtr in order to guarantee the worst-case complexity result in expectation of the algorithm [6]. On the contrary, irerm is applicable to the general problem (1.1) and adjusts the accuracy variables according to the probabilistic requirements (2.1).
A further relevant difference between irerm and sirtr lies in the presence of the parameter $\underline{\theta}$ in irerm, which allows us to avoid the strong assumption, made in sirtr, that every subsampled function evaluation is uniformly bounded from below and above on a set containing the generated sequence $\{x_{k}\}$. As a result, the predicted reduction $\operatorname{Pred}_{k}(\theta)$ and actual reduction $\operatorname{Ared}_{k}(x_{k}+p_{k},\theta)$ have been also modified, by inserting an additional function evaluation $\bar{f}_{k}^{*}$ that is not present in sirtr and ensures that $\theta_{k+1}^{t}\geq\underline{\theta}$ whenever the function and gradient evaluations are sufficiently accurate (see the upcoming Lemma 4.4).
4. Analysis of the proposed algorithm
In this section, we analyze the theoretical properties of the stochastic process underlying Algorithm irerm. Following [15], we denote: $X_{k}$ the random iterate with realization $x_{k}$; $\tilde{Y}_{k}$, the random variable with realization $\tilde{y}_{k}$; $G_{k}$ the random gradient with realization $g_{k}$; $\Delta_{k}$ the random trust-region radius with realization $\delta_{k}$; $P_{k}$ the random step with realization $p_{k}$; $\bar{F}_{k}^{\dagger}$, $\bar{F}_{k}^{*}$, $\bar{F}_{k}^{p}$ the random estimates of $f$ with realizations $\bar{f}_{k}^{\dagger}$, $\bar{f}_{k}^{*}$, $\bar{f}_{k}^{p}$. We denote with $\mathbb{P}_{k-1}(\cdot)$ and $\mathbb{E}_{k-1}[\cdot]$ the probability and expected value conditioned to the $\sigma$-algebra generated up to the beginning of iteration $k$, i.e., the $\sigma$-algebra generated by $\bar{F}_{0}^{\dagger},\bar{F}_{0}^{*},\bar{F}_{0}^{p},\ldots,\bar{F}_{k-1}^{\dagger},\bar{F}_{k-1}^{*},\bar{F}_{k-1}^{p}$, and $G_{0},\ldots,G_{k-1}$, and with $\mathbb{P}_{k-1/2}(\cdot)$ the probability conditioned to the $\sigma$-algebra generated by $\bar{F}_{0}^{\dagger},\bar{F}_{0}^{*},\bar{F}_{0}^{p},\ldots,\bar{F}_{k-1}^{\dagger},\bar{F}_{k-1}^{*},\bar{F}_{k-1}^{p}$, and $G_{0},\ldots,G_{k}$.
In order to show that our algorithm can reach a desired accuracy in the value of the true gradient of $f$ we give an upper bound on the expected value of the hitting time defined below.
Definition 4.1.
Given $\epsilon>0$, the hitting time $\mathcal{K}_{\epsilon}$ is the random variable
$\mathcal{K}_{\epsilon}=\min\{k\geq 0:\ \|\nabla f(X_{k})\|\leq\epsilon\},$
i.e., $\mathcal{K}_{\epsilon}$ is the first iteration such that $\|\nabla f(X_{\mathcal{K}_{\epsilon}})\|\leq\epsilon$.
We proceed as follows. First, we show some properties of the sequence $\{\theta_{k}\}$; second, we introduce conditions which enforce successful iterations; third, we provide results that allow us to rely on the theory given in [9] and to derive the iteration complexity results.
4.1. On the sequence $\{\theta_{k}\}$
We study the properties of the sequence $\{\theta_{k}\}$ that will be crucial in our analysis.
Lemma 4.2.
Suppose Assumption 2.1 holds. The following facts hold true.
  • (i)
    The sequence $\{\theta_{k}\}$ is positive, nonincreasing, and $\theta_{k+1}^{t}\leq\theta_{k}$, for all $k\geq 0$.
  • (ii)
    It holds
    (4.1) $\operatorname{Pred}_{k}(\theta_{k+1}^{t})\geq\theta_{k+1}^{t}\delta_{k}\|g_{k}\|,\quad\forall\ k\geq 0.$
  • (iii)
    The sequence $\{\theta_{k}\}$ is bounded away from zero with $\theta_{k+1}\geq\underline{\theta}>0$, for all $k\geq 0$.
Proof.
(i) We note that $\theta_{0}>0$, and proceed by induction assuming that $\theta_{k}>0$ for some $k>0$. First, suppose that $h(y_{k})=0$. By (3.1), we have that $h(y_{k+1}^{t})=h(\widetilde{y}_{k+1})=0$, which by Assumption 2.1(i) yields $\bar{f}_{k}^{*}=\bar{f}_{k}^{\,\dagger}=f(x_{k})$. Thus, $\operatorname{Pred}_{k}(\theta_{k})={\theta_{k}}\delta_{k}\|g_{k}\|$, the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and $\theta_{k+1}^{t}$ is positive by induction. Now suppose that $h(y_{k})>0$. If inequality $\operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|$ holds, then the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and thus $\theta_{k+1}^{t}>0$ by induction. Otherwise, it must be
(4.2) $\theta_{k}\left(\bar{f}_{k}^{*}-\bar{f}_{k}^{\,\dagger}-(h(y_{k})-h(\widetilde{y}_{k+1}))\right)<-(h(y_{k})-h(\widetilde{y}_{k+1}))<0,$
where the rightmost inequality follows from (3.1). Since $\theta_{k}>0$, (4.2) implies
(4.3) $0<\theta_{k+1}^{*}:=\frac{h(y_{k})-h(\widetilde{y}_{k+1})}{\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}+h(y_{k})-h(\widetilde{y}_{k+1})}<\theta_{k}.$
By the update rule (3.2), we have $\theta_{k+1}^{t}=\theta_{k+1}^{*}$, which is positive from (4.3).
Since Step 4 assigns either $\theta_{k+1}=\theta_{k+1}^{t}$ or $\theta_{k+1}=\theta_{k}$, we conclude that $\theta_{k+1}$ is positive in both cases. Hence, the sequence $\{\theta_{k}\}$ is positive. Moreover, from (3.2) and (4.3) we have $\theta_{k+1}^{t}\leq\theta_{k}$, and $\{\theta_{k}\}_{k\in\mathbb{N}}$ is nonincreasing by Step 4.
(ii) If inequality $\operatorname{Pred}_{k}(\theta_{k})\geq\theta_{k}\delta_{k}\|g_{k}\|$ holds, the update rule (3.2) gives $\theta_{k+1}^{t}=\theta_{k}$, and thus (4.1) trivially holds. Otherwise, $\theta_{k+1}^{t}=\theta_{k+1}^{*}$, and from (4.2) and (4.3) it follows that $\operatorname{Pred}_{k}(\theta)\geq{\theta}\delta_{k}\|g_{k}\|$ if and only if the value $\theta\leq\theta_{k+1}^{*}$ with $\theta_{k+1}^{*}$ given in (4.3). Thus (4.1) is satisfied.
(iii) Note that $\theta_{0}>\underline{\theta}$ by the initial setting and suppose by induction that $\theta_{k}\geq\underline{\theta}$. By (3.3) and (3.4), it holds that $\theta_{k+1}=\theta_{k+1}^{t}$ for successful iterations, $\theta_{k+1}=\theta_{k}$ otherwise. In the former case, we have $\theta_{k+1}\geq\underline{\theta}$ due to the definition of successful iteration, in the latter case $\theta_{k+1}\geq\underline{\theta}$ by induction. Thus, the claim follows. ∎
4.2. Successful iterations
In this section, we introduce conditions that enforce successful iterations, i.e., conditions which guarantee (3.8)–(3.10). The key issue is the characterization of sufficient accuracy on $f$ and $\nabla f$; in the following definition, we formalize such accuracy requirements and call true an iteration where they hold, false otherwise.
Definition 4.3.
For a given iteration $k\geq 0$, let $x_{k},g_{k},\delta_{k},x_{k}+p_{k},y_{k},y_{k+1}^{t},\widetilde{y}_{k+1},y_{k+1}^{g}$ be as in Algorithm irerm. Then, given $\kappa>0$, we say that iteration $k$ is true when
(4.4) $|f(x_{k})-\bar{f}_{k}^{*}|$ $\leq\kappa\min\{\delta_{k}^{2},h(y_{k})\},$
$|f(x_{k})-\bar{f}_{k}^{\,\dagger}|$ $\leq\kappa\min\{\delta_{k}^{2},h(y_{k})\},$
and
(4.5) $\|\nabla f(x_{k})-g_{k}\|$ $\leq\kappa\delta_{k},$
(4.6) $|f(x_{k}+p_{k})-\bar{f}_{k}^{p}|$ $\leq\kappa\delta_{k}^{2}.$
Based on Section 2, we remark that true iterations occur with a certain probability if the level of noise is sufficiently small. In the case of finite-sum minimization, iteration $k$ is true in probability provided that the sample sizes used for computing $\bar{f}_{k}^{*},\,\bar{f}_{k}^{\dagger},\,g_{k},\,\bar{f}_{k}^{p}$ satisfy lower bounds of the form (2.5); in the remaining cases of Section 2, iteration $k$ is true in probability for specified numbers of trials employed in the evaluation of $\bar{f}_{k}^{*},\,\bar{f}_{k}^{\dagger},\,g_{k},\,\bar{f}_{k}^{p}$ and complying with inequalities of the form (2.10). By using Taylor expansion and assuming Lipschitz continuity of $\nabla f$, the second condition in (4.4) and (4.5) imply that model $m_{k}(p)=\bar{f}_{k}^{\dagger}+g_{k}^{T}p$ is fully linear in a ball of center $x_{k}$ and radius $\delta_{k}$, according to the definition of fully linear model given e.g. in [15, Definition 3.1].
Our first result concerns condition (3.10).
Lemma 4.4.
Suppose Assumption 2.1 holds. Let $k$ be an iteration such that condition (4.4) is satisfied with $\kappa\leq(1-r-\underline{\theta})/(2\underline{\theta})$. Then, condition (3.10) holds.
Proof.
Consider Step 3 of Algorithm irerm. If $\theta_{k+1}^{t}=\theta_{k}$ in (3.2), Lemma 4.2(iii) establishes that each $\theta_{k}$ is bounded below by $\underline{\theta}$ and then (3.10) holds. Otherwise, by the proof of Lemma 4.2(i), we know that $h(y_{k})>0$ and $\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}+h(y_{k})-h(\widetilde{y}_{k+1})>0$. Then, condition (4.4) yields
$|\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}|=|\bar{f}_{k}^{\,\dagger}-f(x_{k})+f(x_{k})-\bar{f}_{k}^{*}|\leq 2\kappa h(y_{k}),$
and $h(\widetilde{y}_{k+1})\geq 0$ implies
$0<\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}+h(y_{k})-h(\widetilde{y}_{k+1})\leq|\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}|+h(y_{k})\leq(1+2\kappa)h(y_{k}).$
Then, by using (3.2) and (3.1), we get
$\theta_{k+1}^{t}=\frac{h(y_{k})-h(\widetilde{y}_{k+1})}{\bar{f}_{k}^{\,\dagger}-\bar{f}_{k}^{*}+h(y_{k})-h(\widetilde{y}_{k+1})}\geq\frac{(1-r)h(y_{k})}{(1+2\kappa)h(y_{k})}\geq\underline{\theta},$
where the last inequality follows from the assumptions made on $\kappa$. ∎
The next lemma mimics a standard result holding for deterministic trust-region methods; if the iteration is true and the trust-region is sufficiently small then the iteration is successful. We need the subsequent assumption on $f$.
Assumption 4.5.
The gradient of the function $f$ in (1.1) is Lipschitz continuous with constant $2L$.
Lemma 4.6.
Suppose Assumptions 2.1 and 4.5 hold, and that iteration $k$ is true with $\kappa\leq(1-r-\underline{\theta})/(2\underline{\theta})$. If the trust-region radius $\delta_{k}$ satisfies
(4.7) $\delta_{k}\leq\min\left\{\frac{1}{\eta_{2}}\|g_{k}\|,\ \frac{\underline{\theta}(1-\eta_{1})}{\theta_{0}(3\kappa+L)+\mu}\|g_{k}\|\right\},$
then the iteration is successful.
Proof.
Iteration $k$ is successful when conditions (3.8)–(3.10) are satisfied. Since $k$ is true, Lemma 4.4 guarantees the fulfillment of (3.10), thus we need to show that (3.8) and (3.9) hold when $\delta_{k}$ satisfies (4.7).
Trivially (4.7) implies (3.9). Concerning (3.8), we combine the definition of $\operatorname{Pred}_{k}$ in (3.5), the definition of $\operatorname{Ared}_{k}$ in (3.7), inequality (4.1), condition (3.1) and $\theta_{k+1}^{t}\leq\theta_{k}\leq\theta_{0}$ (see Lemma 4.2(i)), so as to obtain
$\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1}^{t})$ $-\eta_{1}\operatorname{Pred}_{k}(\theta_{k+1}^{t})$
$=(1-\eta_{1})\operatorname{Pred}_{k}(\theta_{k+1}^{t})+\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1}^{t})-\operatorname{Pred}_{k}(\theta_{k+1}^{t})$
$=(1-\eta_{1})\operatorname{Pred}_{k}(\theta_{k+1}^{t})+\theta_{k+1}^{t}(m_{k}(p_{k})-\bar{f}_{k}^{p})$
$\quad+(1-\theta_{k+1}^{t})(h(\widetilde{y}_{k+1})-h(y_{k+1}^{t}))$
$\geq(1-\eta_{1}){\theta_{k+1}^{t}}\delta_{k}\|g_{k}\|+\theta_{k+1}^{t}(m_{k}(p_{k})-\bar{f}_{k}^{p})$
$\quad-(1-\theta_{k+1}^{t})h(y_{k+1}^{t})$
$\geq(1-\eta_{1}){\theta_{k+1}^{t}}\delta_{k}\|g_{k}\|-\theta_{k+1}^{t}|m_{k}(p_{k})-\bar{f}_{k}^{p}|$
$\quad-h(y_{k+1}^{t})$
(4.8) $\geq(1-\eta_{1}){\theta_{k+1}^{t}}\delta_{k}\|g_{k}\|-\theta_{k+1}^{t}|m_{k}(p_{k})-\bar{f}_{k}^{p}|-\mu\delta_{k}^{2}.$
The occurrence that $k$ is true and Assumption 4.5 give
$\theta_{k+1}^{t}|m_{k}(p_{k})$ $-\bar{f}_{k}^{p}|$
$=\theta_{k+1}^{t}\left|\bar{f}_{k}^{\,\dagger}-f(x_{k})+f(x_{k}+p_{k})-\bar{f}_{k}^{p}+(g_{k}-\nabla f(x_{k}))^{T}p_{k}\right.$
$\hskip 35.0pt\left.+f(x_{k})-f(x_{k}+p_{k})+\nabla f(x_{k})^{T}p_{k}\right|$
$\leq\theta_{0}\left(|\bar{f}_{k}^{\,\dagger}-f(x_{k})|+|f(x_{k}+p_{k})-\bar{f}_{k}^{p}|\right)$
[…]
(4.9) $\leq\theta_{0}\left(3\kappa+L\right)\delta_{k}^{2}.$
By plugging (4.9) and (3.10) in (4.8), it follows
$\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1})-\eta_{1}\operatorname{Pred}_{k}(\theta_{k+1})$ $\geq\left(\underline{\theta}(1-\eta_{1})\|g_{k}\|-\left(\theta_{0}\left(3\kappa+L\right)+\mu\right)\delta_{k}\right)\delta_{k},$
and (4.7) guarantees that the acceptance condition (3.8) is satisfied, since the right-hand side of the inequality above is non-negative. ∎
We conclude this section by providing a straightforward lower bound on $\operatorname{Ared}_{k}(x_{k+1},\theta_{k+1})$ in case $k$ is a successful iteration.
Lemma 4.7.
Suppose that iteration $k$ is successful. Then, we have
$\operatorname{Ared}_{k}(x_{k+1},\theta_{k+1})=\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1})\geq\eta_{1}\eta_{2}\theta_{k+1}\delta_{k}^{2}.$
Proof.
If the iteration $k$ is successful, then $x_{k+1}=x_{k}+p_{k}$ and $\theta_{k+1}=\theta_{k+1}^{t}$ according to (3.3). By combining (4.1) with the acceptance conditions (3.8) and (3.9), it follows that
(4.10) $\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1})$$\geq$ $\eta_{1}{\theta_{k+1}}\delta_{k}\|g_{k}\|$
(4.11) $\geq$ $\eta_{1}\eta_{2}\theta_{k+1}\delta_{k}^{2}.$
4.3. Iteration complexity
We conclude our analysis by deriving the expected number $\mathbb{E}[{\mathcal{K}}_{\epsilon}]$ for the hitting time given in Definition 4.1. We rely on results for random processes introduced in [9, §2] and, to this end, we need some technical results that are proved in this section. In particular, we exploit the Lyapunov function:
(4.12) $\psi(x,y,\theta,\delta)=v(\theta{f(x)}+(1-\theta)h(y)+\theta\Sigma)+(1-v)\delta^{2},$
and consider the sequence $\{\psi_{k}\}$ where
(4.13) $\psi_{k}=\psi(x_{k},y_{k},\theta_{k},\delta_{k}),\quad\forall\ k\geq 0.$
Here $v\in(0,1)$ is a prefixed constant to be later specified and $\Sigma$ is a constant such that
(4.14) ${f(x_{k})}-h(y_{k})+\Sigma\geq 0,\quad\forall\ k\geq 0.$
By Assumption 2.1, such a constant exists since $f(x)$ is bounded from below and the values of $h(y_{k})$ are upper bounded by $h_{up}$; one possible choice is $\Sigma=-f_{low}+h_{up}$.
Further, we introduce four events that characterize true iterations.
Definition 4.8.
Given $k\geq 0$, let $\mathcal{G}_{k,1},\mathcal{G}_{k,2},\mathcal{G}_{k,3},\mathcal{G}_{k,4}$ be the events
(4.15) $\mathcal{G}_{k,1}$ =$=\left\{|f(X_{k})-\bar{F}_{k}^{\dagger}|\leq\kappa\min\{\Delta_{k}^{2},h(Y_{k})\}\right\},$
(4.16) $\mathcal{G}_{k,2}$ =$=\left\{|f(X_{k})-\bar{F}_{k}^{*}|\leq\kappa\min\{\Delta_{k}^{2},h(Y_{k})\}\right\},$
(4.17) $\mathcal{G}_{k,3}$ =$=\left\{\|\nabla f(X_{k})-G_{k}\|\leq\kappa\Delta_{k}\right\},$
(4.18) $\mathcal{G}_{k,4}$ =$=\left\{|f(X_{k}+P_{k})-\bar{F}_{k}^{p}|\leq\kappa\Delta_{k}^{2}\right\},$
and $\mathcal{I}_{k},\mathcal{J}_{k}$ the indicator functions
${\mathcal{I}}_{k}={\mathbbm{1}}({\mathcal{G}_{k,1}\cap\mathcal{G}_{k,2}\cap\mathcal{G}_{k,3}}),\quad{\mathcal{J}}_{k}={\mathbbm{1}}({\mathcal{G}_{k,4}}).$
We note that iteration $k$ is true by Definition 4.3 when $\mathcal{I}_{k}\mathcal{J}_{k}=1$.
Assumption 4.9.
Given $k\geq 0$, the events $\mathcal{G}_{k,1}$, $\mathcal{G}_{k,2}$, $\mathcal{G}_{k,3}$ are independent.
Note that Assumption 4.9 is satisfied whenever the random variables $\omega^{\dagger}_{y_{k+1}^{t}}$, $\omega^{*}_{y_{k+1}^{t}}$ and $\omega_{y_{k+1}^{t}}^{g}$ that model the noise are independent.
In the next lemma, we state that the iteration $k$ is true in probability, provided that Assumptions 2.1 and 4.9 hold and the parameter $\mu$ is properly chosen.
Lemma 4.10.
Suppose that Assumptions 2.1 and 4.9 hold. Given a probability $\pi\in(0,1]$, let $\bar{\mu}(\pi^{\frac{1}{4}})>0$ be defined as in Assumption 2.1(ii) and assume $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$. Then we have
(4.19) $\mathbb{P}_{k-1}({\mathcal{I}}_{k}{\mathcal{J}}_{k}=1)\geq\pi.$
Proof.
From the update rule (3.1) and condition $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$, we have
$h({y}^{t}_{k+1})\leq\bar{\mu}(\pi^{\frac{1}{4}})\min\left\{\kappa\delta_{k}^{2},\kappa h(y_{k})\right\},\,\,h({y}^{g}_{k+1})\leq\bar{\mu}(\pi^{\frac{1}{4}})(\kappa\delta_{k}).$
Then, by Assumption 2.1(ii), it follows that
$\mathbb{P}_{k-1}(\mathcal{G}_{k,i})\geq\pi^{\frac{1}{4}},\quad 1\leq i\leq 3.$
Since the events $\mathcal{G}_{k,i}$ are independent by Assumption 4.9, the above inequality yields
$\mathbb{P}_{k-1}(\mathcal{I}_{k}=1)\geq\pi^{\frac{3}{4}}.$
Analogously, from the update rule (3.1) and Assumption 2.1(ii), we have
$\mathbb{P}_{k-1/2}(\mathcal{J}_{k}=1)\geq\pi^{\frac{1}{4}}.$
Then, the thesis follows as in [9, p. 100]. ∎
In the rest of the analysis, we will make use of a series of parameters introduced below which are positive by construction.
Definition 4.11.
Let $\kappa$, $\zeta$ and $v\in(0,1)$ be some scalars satisfying
(4.20) $0<\kappa<\min\left\{\frac{1-r-\underline{\theta}}{2\underline{\theta}},\frac{\eta_{1}\eta_{2}}{2}\right\},$
(4.21) $\zeta>\kappa+\max\left\{\eta_{2},\frac{\theta_{0}(3\kappa+L){+\kappa\bar{\mu}_{up}}}{\underline{\theta}(1-\eta_{1})},\frac{\kappa(2+\eta_{1})}{\eta_{1}}\right\}.$
Let $C_{1},\,C_{2},\,C_{3},\,C_{4}$ and $v$ be the positive constants
(4.22) $\begin{cases}C_{1}=\underline{\theta}\left(\eta_{1}\left(1-\frac{\kappa}{\zeta}\right)-\frac{2\kappa}{\zeta}\right),\\ C_{2}=1+\displaystyle\frac{L+\kappa\bar{\mu}_{up}}{\zeta},\\ C_{3}=\underline{\theta}(\eta_{1}\eta_{2}-2\kappa),\\ C_{4}=L+\zeta+\kappa\bar{\mu}_{up},\end{cases}$
(4.23) $\frac{v}{1-v}=\max\left\{\frac{4\gamma^{2}}{\zeta C_{1}},\frac{2\gamma^{2}}{C_{3}}\right\}.$
We note that the constants $C_{1},\,C_{3}$ are positive due to (4.21) and (4.20), respectively.
The following theorem shows a bound on the expected value $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}]$ from which one easily derives that $\Delta_{k}$ tends to zero almost surely.
Theorem 4.12.
Let $\kappa,\,\zeta,\,C_{1},\,C_{2},\,C_{3},\,C_{4}$ and $v$ as in Definition 4.11. Let $\pi\in\left(\frac{1}{2},1\right)$ be a probability satisfying
(4.24) $\frac{\pi-\frac{1}{2}}{1-\pi}\geq\frac{C_{2}}{C_{1}},\quad 1-\pi\leq\frac{\gamma^{2}-1}{2(\gamma^{4}-1)+2\gamma^{2}C_{4}\max\left\{\frac{4\gamma^{2}}{\zeta C_{1}},\frac{2\gamma^{2}}{C_{3}}\right\}}.$
Suppose that Assumptions 2.1, 4.5 hold. Let $\bar{\mu}(\pi^{\frac{1}{4}})>0$ be defined as in Assumption 2.1(ii) and assume $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$. Let $\psi_{k}$ be defined as in (4.13) and (4.14). Then, there exists $\sigma>0$ such that
(4.25) ${\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}]\leq-\sigma\Delta_{k}^{2},\quad\forall\ k\geq 0.}$
Furthermore,
(4.26) $\sum_{k=0}^{\infty}\Delta_{k}^{2}<\infty$
almost surely.
Proof.
See the Appendix. ∎
We observe that (4.26) implies that $\Delta_{k}$ tends to zero almost surely. Thus, the accuracy requirements given in Definition 4.3 for true iterations are expected to become more stringent as the iterations proceed.
In order to obtain the expected value for the hitting time $\mathcal{K}_{\epsilon}$ given in Definition 4.1, we need the value $\delta^{\dagger}$, independent of $k$, such that if $\delta_{k}<\delta^{\dagger}$, iteration $k$ is true and $k<\mathcal{K}_{\epsilon}$, then the iteration is successful.
Lemma 4.13.
Let Assumptions 2.1, 4.5 hold. Suppose that $\|\nabla f(x_{k})\|>\epsilon$ for some $\epsilon>0$, the iteration $k$ is true, and
(4.27) $\delta_{k}<\delta^{\dagger}:=\min\left\{\frac{\epsilon}{2\kappa},\frac{\epsilon}{2\eta_{2}},\frac{\underline{\theta}\epsilon(1-\eta_{1})}{2\theta_{0}(3\kappa+L)}\right\}.$
Then, iteration k is successful.
Proof.
Conditions (4.5) and (4.27) give $\|g_{k}-\nabla f(x_{k})\|\leq\kappa\delta_{k}<\frac{\epsilon}{2}$. Consequently $\|\nabla f(x_{k})\|>\epsilon$ implies $\|g_{k}\|\geq\frac{\epsilon}{2}$, and iteration $k$ is successful by Lemma 4.6. ∎
Let ${\mathcal{G}}_{k}=\mathcal{G}_{k,1}\cap\mathcal{G}_{k,2}\cap\mathcal{G}_{k,3}\cap\mathcal{G}_{k,4}$ with $\mathcal{G}_{k,1},\mathcal{G}_{k,2},\mathcal{G}_{k,3},\mathcal{G}_{k,4}$ given in Definition 4.8, ${\mathbbm{1}}({\mathcal{G}}_{k})$ be the value of the indicator function for the event ${\mathcal{G}}_{k}$ and $W_{k}$ be the random variable defined as
(4.28) $\begin{cases}W_{0}=1\\ W_{k+1}=2\left({\mathbbm{1}}({\mathcal{G}}_{k})-\frac{1}{2}\right),\quad k=0,1,\ldots\end{cases}$
Clearly, $W_{k}$ takes values ±$\pm 1$. Then, we can prove the following result.
Lemma 4.14.
Let the assumptions of Theorem 4.12 hold, $\delta^{\dagger}$ as in (4.27) and $\mathcal{K}_{\epsilon}$ as in Definition 4.1. Suppose there exists some $j_{\max}\geq 0$ such that $\delta_{\max}=\gamma^{j_{\max}}\delta_{0}$, and $\delta_{0}>\delta^{\dagger}$. Then,
  • i)
    there exists $\lambda>0$ such that $\Delta_{k}\leq\delta_{0}e^{\lambda\cdot j_{\max}}$ for all $k\geq 0$;
  • ii)
    there exists a constant $\delta_{\epsilon}=\delta_{0}e^{\lambda\cdot j_{\epsilon}}$ for some $j_{\epsilon}\leq 0$ such that, for all $k\geq 0$,
    (4.29) $\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}\Delta_{k+1}\geq\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}\min\{\Delta_{k}e^{\lambda W_{k+1}},\delta_{\epsilon}\}.$
  • iii)
    there exists a nondecreasing function :[0,)($\ell:[0,\infty)\rightarrow(0,\infty)$ and a constant $\Theta>0$ such that, for all $k\geq 0$,
    (4.30) $\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}\mathbb{E}_{k-1}[\Psi_{k+1}]\leq\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}\Psi_{k}-\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}\Theta\ell(\Delta_{k}).$
Proof.
The proof parallels that of [9, Lemma 7].
i) Since $\delta_{\max}=\gamma^{j_{\max}}\delta_{0}$, we can set $\lambda=\log(\gamma)>0$, and the thesis follows from imposing $\delta_{k}\leq\delta_{\max}$, for all $k\geq 0$, see Step 4 of Algorithm irerm.
ii) Let us set
(4.31) ${\delta_{\epsilon}=\frac{\epsilon}{\xi}},\quad\text{where }\xi\geq\max\left\{2\kappa,2\eta_{2},\frac{2\theta_{0}(3\kappa+L)}{\underline{\theta}(1-\eta_{1})}\right\},$
and assume that $\delta_{\epsilon}=\gamma^{j_{\epsilon}}\delta_{0}$, for some integer $j_{\epsilon}\leq 0$; notice that we can always choose $\xi$ sufficiently large so that this is true. As a consequence, $\Delta_{k}=\gamma^{i_{k}}\delta_{\epsilon}$ for some integer $i_{k}$.
When $\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}=0$, inequality (4.29) trivially holds. Otherwise, conditioning on $\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}=1$, we can prove that
(4.32) $\Delta_{k+1}\geq\min\{\delta_{\epsilon},\min\{\delta_{\max},\gamma\Delta_{k}\}{\mathbbm{1}}({\mathcal{G}}_{k})+\gamma^{-1}\Delta_{k}(1-{\mathbbm{1}}({\mathcal{G}}_{k}))\}.$
Indeed, taking into account the above equation $\Delta_{k}=\gamma^{i_{k}}\delta_{\epsilon}$, for any realization such that $\delta_{k}>\delta_{\epsilon}$ we have $\delta_{k}\geq\gamma\delta_{\epsilon}$, and the updating rule for $\delta_{k}$ in Step 4 of Algorithm irerm implies $\delta_{k+1}\geq\delta_{\epsilon}$. Now let us consider a realization such that $\delta_{k}\leq\delta_{\epsilon}$. Since $\mathcal{K}_{\epsilon}>k$ and $\delta_{\epsilon}\leq\delta^{\dagger}$, if ${\mathbbm{1}}({\mathcal{G}}_{k})=1$ then we can apply Lemma 4.13 and conclude that $k$ is successful. Hence, by Step 4, we have $\delta_{k+1}=\min\{\delta_{\max},\gamma\delta_{k}\}$. If ${\mathbbm{1}}({\mathcal{G}}_{k})=0$, then we cannot guarantee that $k$ is successful; however, again using the updating rule for $\delta_{k}$ in Step 4 of Algorithm irerm, we can write $\delta_{k+1}\geq\gamma^{-1}\delta_{k}$. Combining these two cases, we get (4.32). If we observe that $\delta_{\max}=\gamma^{j_{\max}}\delta_{0}\geq{\gamma^{j_{\epsilon}}\delta_{0}=\delta_{\epsilon}}$, and recall the definition of $W_{k}$ in (4.28), then equation (4.32) easily yields (4.29).
(iii) The thesis trivially follows from (4.25) with $\ell(\Delta)=\Delta^{2}$ and $\Theta=\sigma$. ∎
We now state the main result.
Theorem 4.15.
Under the assumptions of Lemma 4.14, we have
(4.33) ${\mathbb{E}[\mathcal{K}_{\epsilon}]\leq\frac{{\pi}}{2{\pi}-1}\cdot\frac{\psi_{0}\xi^{2}}{{\sigma}\epsilon^{2}}}+1.$
where $\sigma$ is the scalar in (4.25)
Proof.
The claim coincides with [9, Theorem 2], which holds for any random process $\{(\Phi_{k},\Delta_{k},W_{k})\}_{k\in\mathbb{N}}$ satisfying [9, Assumption 2.1], i.e., conditions $i),ii),iii)$ stated in Lemma 4.14. In particular, following [9], let $A_{n}$ count the iterations for which $\Delta_{k}\geq\delta_{\epsilon}$ on $\mathcal{K}_{\epsilon}>k$, and let $N(k)=\max\{n:A_{n}\leq k\}$. Lemma 2 in [9] shows that $\mathbb{E}[N(\mathcal{K}_{\epsilon}-1)+1]\leq\frac{\psi_{0}\xi^{2}}{\sigma\epsilon^{2}}$. Then, the Wald’s Identity [9, Theorem 1] gives $\mathbb{E}[A_{N(\mathcal{K}_{\epsilon}-1)+1}]=\frac{\pi}{2\pi-1}\mathbb{E}[N(\mathcal{K}_{\epsilon}-1)+1]$, and the claim follows by $A_{N(\mathcal{K}_{\epsilon}-1)+1}\geq\mathcal{K}_{\epsilon}-1$ and the technical Lemma 1 in [9]. ∎
An iteration complexity result analogous to Theorem 4.15 can be found in [4, Theorem 3] for the algorithm sirtr applicable to finite-sum minimization and based on the combination of the Inexact Restoration framework and the stochastic first-order trust-region method. The probabilistic requirements in [4] are imposed only on the stochastic gradients and involve bounds of order $\delta_{k}$, similarly to (4.5), whereas irerm requires the more stringent conditions (4.4) and (4.6). On the other hand, the convergence analysis in [4] is carried out under the assumption that the objective function $f$ is bounded from above on a set containing the sequence of the iterates, which is not required in our analysis.
5. Numerical experiments
In this section, we present results from our numerical validation of irerm. Following [15, §6.1], we address the solution of the least-squares problem
$\min_{x\in\mathbb{R}^{n}}f(x)=\sum_{i=1}^{m}(f_{i}(x))^{2},$
where each $f_{i}$, $i=1,\ldots,m$, is a smooth function and suppose that the function evaluations are corrupted by multiplicative noise, i.e.,
(5.1) $f(x,\xi)=\sum_{i=1}^{m}((1+\xi_{i})f_{i}(x))^{2},$
with $\xi_{i}$, $i=1,\ldots,m$, generated from a uniform random distribution on $[-\sigma,\sigma]$ for some $\sigma>0$. We consider 17 nonlinear least-squares problems taken from [20], which represent the overall set of problems extracted from References n. $8,21,26,36$ in [20, Section 2]. The selected problems are reported in Table 1. Note that the dimension of the unknown is set to $n=100$ for each problem.
Since $f$ has the form (2.6), i.e., $f(x)=\mathbb{E}_{\xi}[f(x,\xi)]$, we apply irerm as discussed in Section 2. Thus we set the constraint function in (1.2) equal to $h(y)=\sqrt{y}$, where $p(y)=\lceil 1/y\rceil\in\mathbb{N}$ represents the number of samples for a given $y\in(0,1]$, and evaluate function and derivative approximations as in (2.7) and (2.8), respectively.
We apply irerm and a first-order version of storm [15] to the problems in Table 1 and compare the obtained results. Our implementation of the two algorithms is as follows. Firstly, irerm and storm are implemented on the basis of the theoretical prescriptions on function and derivative approximations derived in this work and in [15], respectively; the resulting algorithms are denoted as irerm_v1 and storm_v1. Secondly, irerm and storm are implemented ignoring the theoretical prescriptions and following Algorithm 3 in [15, p. 485]; the resulting algorithms are denoted as irerm_v2 and storm_v2.
Specifically, in irerm_v1 we set the number of samples as
$p(y_{k+1}^{t})=\left\lceil\frac{1}{\mu^{2}\min\{h(y_{k})^{2},\delta_{k}^{4}\}}\right\rceil,\quad p(y_{k+1}^{g})=\left\lceil\frac{1}{\mu^{2}\delta_{k}^{2}}\right\rceil,$
with $\mu=0.99$, so that the inequalities (3.1) at Step 1 of Algorithm irerm are satisfied. As for storm_v1, we implement Algorithm 1 in [15, p. 454] by computing the average function evaluations with sample size $\left\lceil\frac{1}{\delta_{k}^{4}}\right\rceil$ and the average gradient evaluation with sample size $\left\lceil\frac{1}{\delta_{k}^{2}}\right\rceil$, according to the discussion in [15, §5].
On the other hand, following Algorithm 3 in [15, p. 485], we adopt $\max\left\{10+k,\left\lceil\frac{1}{\delta_{k}^{2}}\right\rceil\right\}$ as the heuristic choice for the number of samples in storm_v2, and employ the same value for $p(y_{k+1}^{t})$ and $p(y_{k+1}^{g})$ in irerm_v2. Such a choice for the sample sizes allows for the worsening of the infeasibility measure, in the spirit of the IR framework, although it does not strictly comply with Step 1 of Algorithm irerm.
In both versions of irerm, the parameters are set as $y_{0}=1$, $\theta_{0}=0.9$, $\underline{\theta}=10^{-8}$, $\delta_{0}=1$, $\delta_{\max}=10$, $\gamma=2$, $\eta=0.1$, $\eta_{2}=10^{-3}$. Likewise, both versions of storm are implemented with $\delta_{0}=1$, $\delta_{\max}=10$, $\gamma=2$, $\eta=0.1$, $\eta_{2}=10^{-3}$. Hence, all algorithms are equipped with the same parameter values, except for the additional parameters $y_{0},\theta_{0},\underline{\theta}$ required in irerm. Note that the parameter values common to irerm and storm are the same chosen in the numerical experiments of [15, §6].
For each problem and algorithm, we performed $10$ runs using the level of noise $\sigma=10^{-1}$ as in [15, §6] and the initial guesses given in [20]. We adopt the number of samples employed in the average function and gradient approximations as the measure of the cost of the algorithms; more precisely, we use a counter Cost that is increased, at each function or gradient evaluation, by the number of samples employed at such evaluation. Furthermore, we impose a computational budget equal to $10^{5}(n+1)$ evaluations for irerm_v1 and storm_v1, and equal to $10^{4}(n+1)$ evaluations for irerm_v2 and storm_v2. We observe that irerm requires three function evaluations and one gradient evaluation, whereas storm asks for two function evaluations and one gradient evaluation. For all algorithms, we stop the iterations when either the maximum computational budget is exceeded or a maximum number of $k_{\max}=500$ iterations is met. Note that the upper bound $k_{\max}$ was never reached in our numerical tests.
Problem Problem name Source $n$ $m$
p1 Chained Rosenbrock function [20, Problem 2.1] 100 $2(n-1)$
p2 Chained Wood function [20, Problem 2.2] 100 $3(n-2)$
p3 Chained Powell singular function [20, Problem 2.3] 100 $2(n-2)$
p4 Chained Gragg and Levy function [20, Problem 2.4] 100 $25(n-2)/$ 2
p5 Generalized Broyden tridiagonal function [20, Problem 2.5] 100 $n$
p6 Generalized Broyden banded function [20, Problem 2.6] 100 $n$
p7 Chained Freudstein and Roth function [20, Problem 2.7] 100 $2(n-1)$
p8 Toint quadratic merging problem [20, Problem 2.9] 100 $3(n-2)$
p9 Chained exponential problem [20, Problem 2.10] 100 $2(n-1)$
p10 [20, Problem 2.53] 100 $n$
p11 [20, Problem 2.71] 100 $n$
p12 [20, Problem 2.72] 100 $n$
p13 nondquar [20, Problem 2.79] 100 $n$
p14 sinquad [20, Problem 2.81] 100 $n$
p15 edensch [20, Problem 2.82] 100 $3(n-1)$
p16 genhumps [20, Problem 2.83] 100 $3(n-1)$
p17 errinros (modified) [20, Problem 2.84] 100 $2(n-1)$
Table 1. Test problems description.
We now summarize the numerical results reported in this section. In Table 2, we show the lowest values of $f$ achieved at termination over the $10$ runs by algorithms irerm_v1, storm_v1, irerm_v2, storm_v2, whereas in Table 3 we report the average function value of $f$ and the standard deviation reached at termination over the 10 runs by all algorithms, respectively. Focusing on problems where both versions v1 and v2 of algorithms irerm and storm show significant differences at termination, in Figure 1 (resp. Figure 2) we plot the best runs of irerm_v1 and storm_v1 (resp. irerm_v2 and storm_v2), i.e., the runs where the smallest value of the exact function $f$ is achieved at termination. For each problem, we display the value of the exact function $f$ versus the computational cost (Cost), and the trust-region radius $\delta_{k}$ versus the computational cost (Cost).
Problem irerm_v1 storm_v1 irerm_v2 storm_v2
p1 4.93e+01 4.90e+01 4.73e+01 4.78e+01
p2 1.94e+02 1.46e+02 1.59e+02 1.84e+02
p3 4.15e--02 3.14e--02 1.17e--04 6.47e--03
p4 1.27e+01 1.27e+01 1.26e+01 1.26e+01
p5 4.18e--05 8.23e--05 8.69e--07 6.29e--06
p6 2.73e--03 4.14e--03 1.84e--07 2.76e--05
p7 6.01e+03 6.01e+03 6.00e+03 6.00e+03
p8 2.17e+02 2.17e+02 2.16e+02 2.17e+02
p9 1.96e+01 1.96e+01 1.91e+01 1.91e+01
p10 2.63e--08 7.12e--08 9.48e--08 1.40e--08
p11 4.68e--06 1.54e--05 2.61e--08 4.83e--07
p12 1.29e--05 1.02e--04 4.54e--07 3.30e--06
p13 2.96e--01 2.23e--01 1.84e--01 1.54e--01
p14 1.16e--01 1.29e--01 7.86e--02 7.06e--02
p15 2.94e+02 2.94e+02 2.94e+02 2.94e+02
p16 1.25e+06 6.74e+05 1.84e+03 3.38e+03
p17 3.92e+01 3.92e+01 3.91e+01 3.90e+01
Table 2. Lowest value of $f$ reached at termination over 10 runs for irerm and storm.
Problem irerm_v1 storm_v1 irerm_v2 storm_v2
p1 4.93e+01$\pm$3.79e--02 4.92e+01$\pm$1.01e--01 4.78e+01$\pm$5.40e--01 4.85e+01$\pm$6.74e--01
p2 1.95e+02$\pm$1.19e+00 1.56e+02$\pm$5.73e+00 1.82e+02$\pm$1.07e+01 1.87e+02$\pm$1.03e+00
p3 9.63e--02$\pm$7.01e--02 7.11e--02$\pm$2.34e--02 2.85e--02$\pm$3.01e--02 8.31e--03$\pm$1.45e--03
p4 1.30e+01$\pm$7.81e--01 1.27e+01$\pm$5.58e--03 1.27e+01$\pm$1.75e--01 1.26e+01$\pm$6.52e--03
p5 2.06e--03$\pm$8.64e--04 1.58e--03$\pm$1.06e--03 1.81e--05$\pm$1.61e--05 2.59e--05$\pm$2.11e--05
p6 6.76e--03$\pm$3.24e--03 1.03e--02$\pm$3.20e--03 5.45e--05$\pm$7.39e--05 1.89e--04$\pm$8.23e--05
p7 6.01e+03$\pm$6.77e--02 6.01e+03$\pm$7.17e--02 6.00e+03$\pm$3.26e+00 6.01e+03$\pm$1.12e+00
p8 4.15e+03$\pm$6.22e+03 2.17e+02$\pm$8.22e--03 2.46e+02$\pm$2.52e+01 2.17e+02$\pm$5.18e--03
p9 1.97e+01$\pm$5.36e--02 1.97e+01$\pm$6.86e--02 1.92e+01$\pm$4.67e--02 1.91e+01$\pm$1.29e--02
p10 2.80e--05$\pm$3.64e--05 1.32e--05$\pm$1.21e--05 9.92e--07$\pm$5.61e--07 1.87e--07$\pm$1.62e--07
p11 6.85e--05$\pm$5.05e--05 3.01e--05$\pm$1.12e--05 1.86e--06$\pm$1.48e--06 9.16e--07$\pm$2.47e--07
p12 1.65e--04$\pm$1.54e--04 3.05e--04$\pm$1.21e--04 8.94e--06$\pm$1.57e--05 8.35e--06$\pm$3.24e--06
p13 3.23e--01$\pm$2.66e--02 3.03e--01$\pm$2.85e--02 1.90e--01$\pm$6.66e--03 1.59e--01$\pm$4.94e--03
p14 1.25e--01$\pm$6.48e--03 1.30e--01$\pm$1.38e--04 8.08e--02$\pm$1.62e--03 7.28e--02$\pm$1.06e--03
p15 2.94e+02$\pm$1.33e--02 2.94e+02$\pm$8.39e--03 2.94e+02$\pm$1.46e--05 2.94e+02$\pm$2.07e--02
p16 1.27e+06$\pm$6.14e+03 1.13e+06$\pm$1.68e+05 2.04e+03$\pm$1.08e+02 3.88e+03$\pm$3.65e+02
p17 5.13e+01$\pm$1.61e+01 3.93e+01$\pm$4.00e--02 3.97e+01$\pm$9.28e--01 3.91e+01$\pm$3.10e--02
Table 3. Average function value $f$ and standard deviation at termination over 10 runs for irerm and storm.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1. Best runs of irerm_v1 and storm_v1 in terms of value of $f$ at termination. From left to right: value of true function $f$ vs computational budget and trust-region radius $\delta_{k}$ vs computational budget. From top to bottom: Problems p5, p10, p11, p12.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2. Best runs of irerm_v2 and storm_v2 in terms of value of $f$ at termination. From left to right: value of true function $f$ vs computational budget and trust-region radius $\delta_{k}$ vs computational budget. From top to bottom: Problems p5, p10, p11, p12.
Comparing the smallest value of $f$ achieved by irerm_v1 and storm_v1, we observe in Figure 1 and Table 2 that the two solvers show comparable performances over the selected problems. With respect to storm_v1, irerm_v1 reaches the smallest value $9$ times out of $17$; notably, the minimum value obtained by irerm is at least half of that of storm on $4$ problems (p5, p10, p11, p12). On the other hand, in the case where storm_v1 reaches the smallest value of $f$ at termination, such value is comparable with the value attained by irerm_v1, except for problem p16. In most iterations, the trust-region radius selected in irerm_v1 is larger than in storm_v1; this fact can be ascribed to the acceptance rule of the iterates, which depends on the decrease of a convex combination of the values of $\bar{f}$ and $h$, instead of using only the values $\bar{f}$ as in storm_v1. We observe that larger trust-region radii generally imply looser accuracy requirements on functions and derivatives. This feature compensates the per iteration extra cost or irerm compared to the per iteration cost of storm and makes irerm competitive with storm. The average value of $f$ reached at termination and the standard deviation of the two algorithms is summarized in Table 3. The average values obtained by irerm_v1 and storm_v1 are similar in most cases, with the exception of problems p6 and p8.
Comparing the smallest value of $f$ achieved by irerm_v2 and storm_v2 displayed in Figure 2 and Table 2, we note that irerm_v2 reaches the smallest value $12$ times out of $17$ and compares well with storm_v2 in the remaining problems. In several iterations, the trust-region radius selected in irerm_v2 is larger than in storm_v2. Table 3 shows that the average values over 10 runs are similar for irerm_v2 and storm_v2 except for problems p6 and p10.
The obtained results confirm that solving an unconstrained noisy problem via a suitable constrained formulation and the Inexact Restoration approach is viable and compares well with storm algorithm. Interestingly, the best results over multiple runs are good even for irerm_v2, which represents an heuristic version of our algorithm where the theoretical prescriptions are ignored. We also note that the proposed algorithm tends to compute larger trust-region radii than those of storm, which leads in some cases to smaller values of $f$ at termination of the best runs.
6. Conclusions and future work
We proposed a trust-region algorithm with first-order random models suitable for the minimization of unconstrained noisy functions. The proposed algorithm, named irerm, is based on a constrained reformulation of the minimization problem typical of the Inexact Restoration approach [1, 12, 14, 17], where satisfying the constraint is equivalent to computing both the function and its gradient exactly. To our knowledge, this is the first attempt to adopt the Inexact Restoration approach for a general problem where the evaluation of both the objective function and its gradient is random and sufficiently accurate in probability, and to provide a theoretical analysis under conditions viable for the context. At each iteration, irerm first enforces the sufficient accuracy in probability of the function and gradient estimates, and then employs an acceptance test involving both the noisy function and the infeasibility measure. Based on the theoretical analysis in [9, 15], we derived the expected number of iterations needed to achieve an approximate first-order optimality point. A preliminary numerical illustration of irerm on a collection of nonlinear least-squares problems is also presented. As future work, we plan to deepen the numerical investigation of irerm as well as design a generalized version of irerm with second-order random models.
7. Appendix
This appendix is devoted to the proof of Theorem 4.12.
Proof.
The proof follows closely [15, Theorem 4.11]. Given $\psi_{k}$ in (4.13), the successive difference $\psi_{k+1}-\psi_{k}$ can be written as
$\psi_{k+1}-\psi_{k}$ $=v(\theta_{k+1}f(x_{k+1})+(1-\theta_{k+1})h(y_{k+1}))$
$\quad-v(\theta_{k}{f(x_{k})}+(1-\theta_{k})h(y_{k}))+v(\theta_{k+1}-\theta_{k})\Sigma+(1-v)(\delta_{k+1}^{2}-\delta_{k}^{2})$
$=v\left(\theta_{k+1}({f(x_{k+1})}-{f(x_{k})})+(1-\theta_{k+1})(h(y_{k+1})-h(y_{k}))\right)$
$\quad+v(\theta_{k+1}-\theta_{k})({f(x_{k})}-h(y_{k})+\Sigma)+(1-v)(\delta_{k+1}^{2}-\delta_{k}^{2})$
$\leq v\left(\theta_{k+1}({f(x_{k+1})}-{f(x_{k})})+(1-\theta_{k+1})(h(y_{k+1})-h(y_{k}))\right)$
(7.1) $\quad+(1-v)(\delta_{k+1}^{2}-\delta_{k}^{2}),$
where the last inequality follows from (4.14) and the monotonicity of $\{\theta_{k}\}$ (see Lemma 4.2(i)). If $k$ is unsuccessful, then $x_{k+1}=x_{k}$, $y_{k+1}=y_{k}$, $\delta_{k+1}=\delta_{k}/\gamma$ and (7) reads
(7.2) $\psi_{k+1}-\psi_{k}\leq(1-v)\left(\frac{1-\gamma^{2}}{\gamma^{2}}\right)\delta_{k}^{2}<0.$
To proceed, we distinguish the case $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$ from the case $\|\nabla f(x_{k})\|<\zeta\delta_{k}$.
(i) Case $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$. We analyse separately the case of true and false iteration $k$. If the iteration $k$ is true, we can combine (4.5) with $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$ to obtain
(7.3) $\|g_{k}\|\geq\|\nabla f(x_{k})\|-\kappa\delta_{k}\geq(\zeta-\kappa)\delta_{k}\geq\max\left\{\eta_{2},\frac{\theta_{0}(3\kappa+L)+\mu}{\underline{\theta}(1-\eta_{1})}\right\}\delta_{k},$
where the last inequality is due to condition (4.21) and the upper bound $\mu\leq\kappa\bar{\mu}_{up}$, which follows from the hypothesis $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$ and Assumption 2.1(ii). Since $k$ is true, $\kappa$ satisfies condition (4.20), and inequality (7.3) holds, Lemma 4.6 implies that $k$ is successful. Therefore, (3.3) and inequality (7) yield
$\psi_{k+1}-\psi_{k}$ $\leq v(\theta_{k+1}(f(x_{k}+p_{k})-f(x_{k}))+(1-\theta_{k+1})(h(y_{k+1})-h(y_{k})))$
$+(1-v)(\gamma^{2}-1)\delta_{k}^{2}.$
By summing and subtracting the quantities $\bar{f}_{k}^{*}$ and $\bar{f}_{k}^{p}$ in the above inequality, we get
$\psi_{k+1}-\psi_{k}$ $\leq v(\theta_{k+1}(\bar{f}_{k}^{p}-\bar{f}_{k}^{*})+(1-\theta_{k+1})(h(y_{k+1})-h(y_{k})))+(1-v)(\gamma^{2}-1)\delta_{k}^{2}$
$\ \ +v\theta_{k+1}(f(x_{k}+p_{k})-\bar{f}_{k}^{p}+\bar{f}_{k}^{*}-f(x_{k}))$
$\leq v(\theta_{k+1}(\bar{f}_{k}^{p}-\bar{f}_{k}^{*})+(1-\theta_{k+1})(h(y_{k+1})-h(y_{k})))+(1-v)(\gamma^{2}-1)\delta_{k}^{2}$
$\ \ +v\theta_{k+1}(|f(x_{k}+p_{k})-\bar{f}_{k}^{p}|+|\bar{f}_{k}^{*}-f(x_{k})|).$
By the definition of $\operatorname{Ared}_{k}$ in (3.7), the fact that $y_{k+1}=y_{k+1}^{t}$ due to $k$ being successful, and (4.4)-(4.6) we get
(7.4) $\psi_{k+1}-\psi_{k}\leq-v\operatorname{Ared}_{k}(x_{k}+p_{k},\theta_{k+1})+(1-v)(\gamma^{2}-1)\delta_{k}^{2}+2v\theta_{k+1}\kappa\delta_{k}^{2}.$
Since we know that (4.10) holds for successful iterations, from the previous inequality we easily get
(7.5) $\psi_{k+1}-\psi_{k}\leq-v\eta_{1}\theta_{k+1}\delta_{k}\|g_{k}\|+(1-v)(\gamma^{2}-1)\delta_{k}^{2}+2v\theta_{k+1}\kappa\delta_{k}^{2}.$
We exploit once again the fact that $k$ is true and $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$, and observe that
$\|g_{k}\|\geq\|\nabla f(x_{k})\|-\kappa\delta_{k}\geq\left(1-\frac{\kappa}{\zeta}\right)\|\nabla f(x_{k})\|,$
and (7.5) yields
$\psi_{k+1}-\psi_{k}$ $\leq-v\eta_{1}\theta_{k+1}\left(1-\frac{\kappa}{\zeta}\right)\|\nabla f(x_{k})\|\delta_{k}+(1-v)(\gamma^{2}-1)\delta_{k}^{2}+\frac{2v\theta_{k+1}\kappa\delta_{k}\|\nabla f(x_{k})\|}{\zeta}$
$=v\theta_{k+1}\left(\frac{2\kappa}{\zeta}-\eta_{1}\left(1-\frac{\kappa}{\zeta}\right)\right)\|\nabla f(x_{k})\|\delta_{k}+(1-v)(\gamma^{2}-1)\delta_{k}^{2}.$
Note that $\frac{2\kappa}{\zeta}-\eta_{1}\left(1-\frac{\kappa}{\zeta}\right)<0$ due to (4.21), and $\theta_{k+1}=\theta_{k+1}^{t}\geq\underline{\theta}$ since the acceptance condition (3.10) holds when $k$ is successful. Then, by the definition of $C_{1}$ in (4.22), it follows
(7.6) $\psi_{k+1}-\psi_{k}\leq-vC_{1}\delta_{k}\|\nabla f(x_{k})\|+(1-v)(\gamma^{2}-1)\delta_{k}^{2},$
and by $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$ and (4.23) we have that
$-vC_{1}\delta_{k}\|\nabla f(x_{k})\|+(1-v)(\gamma^{2}-1)\delta_{k}^{2}\leq(-vC_{1}\zeta+(1-v)(\gamma^{2}-1))\delta_{k}^{2}<0.$
If the iteration $k$ is false, $k$ can be either a successful or an unsuccessful iteration. If $k$ is successful, then $x_{k+1}=x_{k}+p_{k}$, and applying the Descent Lemma for functions with Lipschitz continuous gradient to $f$ leads to the following chain of inequalities
$f(x_{k+1})-f(x_{k})$ $\leq\nabla f(x_{k})^{T}p_{k}+L\|p_{k}\|^{2}$
(7.7) $\leq\|\nabla f(x_{k})\|\delta_{k}+L\delta_{k}^{2}$
(7.8) $\leq\left(1+\frac{L}{\zeta}\right)\|\nabla f(x_{k})\|\delta_{k},$
where the last inequality follows from $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$. Furthermore, we have
(7.9) $h(y_{k+1})-h(y_{k})=h(y_{k+1}^{t})-h(y_{k})\leq\mu\delta_{k}^{2},$
due to (3.1). By plugging (7.8) and (7.9) inside (7), we get
$\psi_{k+1}-\psi_{k}$ $\leq v\left(\theta_{k+1}\left(1+\frac{L}{\zeta}\right)\|\nabla f(x_{k})\|\delta_{k}+(1-\theta_{k+1})\mu\delta_{k}^{2}\right)+(1-v)(\gamma^{2}-1)\delta_{k}^{2}$
(7.10) $\leq vC_{2}\|\nabla f(x_{k})\|\delta_{k}+(1-v)(\gamma^{2}-1)\delta_{k}^{2},$
where the second inequality follows from $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$, the bounds $10<\theta_{k+1}<$ 1, the hypothesis $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})\leq\kappa\bar{\mu}_{up}$ and the definition of $C_{2}$ in (4.22).
If $k$ is unsuccessful, then (7.2) holds. Since the right-hand side of (7) is non-negative, we conclude that inequality (7) holds either if $k$ is successful or unsuccessful.
Let $\Psi_{k}$ be the random variable with realization $\psi_{k}$. We are now ready to consider the expected value of $\Psi_{k+1}-\Psi_{k}$ conditioned to the past and the event $\{\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}\}$. By recalling Lemma 4.10 and combining (7.6) and (7), we obtain
$\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}]$
$\quad\qquad\leq\pi(-vC_{1}\Delta_{k}\|\nabla f(X_{k})\|+(1-v)(\gamma^{2}-1)\Delta_{k}^{2})+$
$\quad\qquad(1-\pi)(vC_{2}\|\nabla f(X_{k})\|\Delta_{k}+(1-v)(\gamma^{2}-1)\Delta_{k}^{2})$
(7.11) $\quad\qquad=v\|\nabla f(X_{k})\|\Delta_{k}(-C_{1}\pi+(1-\pi)C_{2})+(1-v)(\gamma^{2}-1)\Delta_{k}^{2}.$
The choice of $v$ in (4.23) and the choice of $\pi$ in (4.24) imply $\frac{C_{1}}{2}\geq\frac{2\gamma^{2}(1-v)}{v\zeta}$ and $C_{1}\pi-(1-\pi)C_{2}\geq\frac{C_{1}}{2}$, respectively. Consequently
(7.12) $C_{1}\pi-(1-\pi)C_{2}\geq\frac{2(\gamma^{2}-1)(1-v)}{v\zeta}.$
Thus (7.11) leads to
$\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}]$ $\leq-\frac{2(1-v)(\gamma^{2}-1)}{\zeta}\|\nabla f(X_{k})\|\Delta_{k}$
$\qquad+(1-v)(\gamma^{2}-1)\Delta_{k}^{2}$
$\leq-2(1-v)(\gamma^{2}-1)\Delta_{k}^{2}+(1-v)(\gamma^{2}-1)\Delta_{k}^{2}$
$=-(1-v)(\gamma^{2}-1)\Delta_{k}^{2},$
where the third inequality is again due to $\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}$. Thus,
(7.13) $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}]\leq-(1-v)(\gamma^{2}-1)\Delta_{k}^{2}.$
(ii) Case $\|\nabla f(x_{k})\|<\zeta\delta_{k}$. Again, we analyse true and false iterations separately. If the iteration $k$ is true and successful, then, reasoning as in Case (i), (7.4) still holds. Thus, (7.4) and (4.11) yield
$\psi_{k+1}-\psi_{k}$ $\leq-v\eta_{1}\eta_{2}\theta_{k+1}\delta_{k}^{2}+(1-v)(\gamma^{2}-1)\delta_{k}^{2}+2v\theta_{k+1}\kappa\delta_{k}^{2}$
$=v\theta_{k+1}(2\kappa-\eta_{1}\eta_{2})\delta_{k}^{2}+(1-v)(\gamma^{2}-1)\delta_{k}^{2}.$
We observe that $02\kappa-\eta_{1}\eta_{2}<$ 0 by (4.20) and $\theta_{k+1}=\theta_{k+1}^{t}\geq\underline{\theta}$ due to the acceptance condition (3.10) holding for successful iterations. Then, by recalling the definition of $C_{3}$ in (4.22), we get
(7.14) $\psi_{k+1}-\psi_{k}\leq(-vC_{3}+(1-v)(\gamma^{2}-1))\delta_{k}^{2}.$
The form of $v$ in (4.23) gives $vC_{3}\geq 2\gamma^{2}(1-v)$ and consequently $vC_{3}\geq 2(\gamma^{2}-1)(1-v)$, then (7.14) leads to
$\psi_{k+1}-\psi_{k}\leq-(1-v)(\gamma^{2}-1)\delta_{k}^{2}.$
If $k$ is unsuccessful, then the sufficient decrease of $\psi_{k+1}-\psi_{k}$ is guaranteed by (7.2). Since $(\gamma^{2}-1)/\gamma^{2}<\gamma^{2}-1$, we conclude that (7.2) holds for all true iterations.
If iteration $k$ is false and successful, then $x_{k+1}=x_{k}+p_{k}$, (7.7) gives
(7.15) $f(x_{k+1})-f(x_{k})$ $\leq\left(\zeta+L\right)\delta_{k}^{2},$
where the last inequality follows from $\|\nabla f(x_{k})\|<\zeta\delta_{k}$. Therefore, (7), (7.9) and (7.15) yield to
$\psi_{k+1}-\psi_{k}$ $\leq v(\theta_{k+1}(L+\zeta)\delta_{k}^{2}+(1-\theta_{k+1})\mu\delta_{k}^{2})+(1-v)(\gamma^{2}-1)\delta_{k}^{2}$
(7.16) $\leq(vC_{4}+(1-v)(\gamma^{2}-1))\delta_{k}^{2},$
where the second inequality follows from the fact that $0<\theta_{k+1}\leq 1$, the hypothesis $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})\leq\kappa\bar{\mu}_{up}$ and the definition of $C_{4}$ given in (4.22).
If $k$ is unsuccessful, then (7.2) holds, and thus $\psi_{k+1}-\psi_{k}<0$. Since the right-hand side of (7) is non-negative, we conclude that inequality (7) holds for all false iterations.
By taking the expected value of $\Psi_{k+1}-\Psi_{k}$ conditioned to the past and the event $\|\nabla f(X_{k})\|<\zeta\Delta_{k}$, and combining (7.2) with (7), we obtain
$\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|<\zeta\Delta_{k}]$
$\qquad\leq\pi(1-v)\left(\frac{1-\gamma^{2}}{\gamma^{2}}\right)\Delta_{k}^{2}+(1-\pi)(vC_{4}+(1-v)(\gamma^{2}-1))\Delta_{k}^{2}$
$\qquad=(\pi\pm 1)(1-v)\left(\frac{1-\gamma^{2}}{\gamma^{2}}\right)\Delta_{k}^{2}+(1-\pi)(vC_{4}+(1-v)(\gamma^{2}-1))\Delta_{k}^{2}$
$\qquad=(1-v)\left(\frac{1-\gamma^{2}}{\gamma^{2}}\right)\Delta_{k}^{2}+(1-\pi)\left(vC_{4}+(1-v)\left(\frac{\gamma^{4}-1}{\gamma^{2}}\right)\right)\Delta_{k}^{2}.$
Using (4.23), the choice of $\pi$ in (4.24) gives $(1-\pi)\left(vC_{4}+(1-v)\left(\frac{\gamma^{4}-1}{\gamma^{2}}\right)\right)\leq\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)$. Consequently,
(7.17) $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|<\zeta\Delta_{k}]\leq-\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)\Delta_{k}^{2}.$
Finally, by combining (7.13) and (7.17) and observing that $(\gamma^{2}-1)/(2\gamma^{2})<\gamma^{2}-1$, we obtain
(7.18) $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}]\leq-\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)\Delta_{k}^{2},$
and we conclude that (4.25) holds with $\sigma=\frac{1}{2}(1-v)(\gamma^{2}-1)/\gamma^{2}$.
In order to prove (4.26), summing (4.25) over $k=0,\ldots,K$, one gets
$\sum_{k=0}^{K}\sigma\Delta_{k}^{2}\leq\sum_{k=0}^{K}\mathbb{E}_{k-1}[\Psi_{k}-\Psi_{k+1}].$
By taking the total expected value and recalling the definition of conditional expected value, we obtain
$\mathbb{E}\left[\sum_{k=0}^{K}\sigma\Delta_{k}^{2}\right]$ $=\sum_{k=0}^{K}\mathbb{E}\left[\mathbb{E}_{k-1}[\Psi_{k}-\Psi_{k+1}]\right]$
$=\mathbb{E}\left[\sum_{k=0}^{K}(\Psi_{k}-\Psi_{k+1})\right]$
$\leq\mathbb{E}[\Psi_{0}-\Psi_{low}],$
where the last inequality is due to the fact that $\{\Psi_{k}\}$ is bounded from below by a constant $\Psi_{low}$ for all $k\geq 0$, thanks to Assumption 2.1(i) and (iii). By taking the limit for $K\rightarrow\infty$ and applying the monotone convergence theorem on the left-hand side of the inequality, we get $\mathbb{E}\left[\sum_{k=0}^{\infty}\Delta_{k}^{2}\right]<\infty$ from which (4.26) immediately follows. ∎